OpenStudy (anonymous):

Monica spent 3/5 hours listening to tapes of beethoven and brahms. She spent 1/4 listening to beethoven . How many hours were spent listening to bramhs?

jimthompson5910 (jim_thompson5910):

let x = # of hours listening to bramhs so if she spent a total of 3/5 hrs on both and 1/4 hr listening to beethoven, then this means x + 1/4 = 3/5

jimthompson5910 (jim_thompson5910):

solve for x

OpenStudy (anonymous):

how would i solve for x

jimthompson5910 (jim_thompson5910):

by subtracting 1/4 from both sides

OpenStudy (anonymous):

so it would be 2/5 right?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

the denominators are NOT equal, so you can't do what you just did

OpenStudy (anonymous):

so then what would i do can u show me step by step im not understanding this

jimthompson5910 (jim_thompson5910):

x + 1/4 = 3/5 x = 3/5 - 1/4 x = 12/20 - 5/20 ... notice how I'm making the denominators equal to the LCD x = ??

OpenStudy (anonymous):

is it 11/20

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

7/20

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

thanks for having patience with me

jimthompson5910 (jim_thompson5910):

you're welcome